OpenStudy (kitkat1):

Medal for best answer!!! What is the exact value of the expression the square root of 72 − the square root of 8 + the square root of 128 ? Simplify if possible. 8the square root of 2 12the square root of 2 8the square root of 3 12the square root of 3

OpenStudy (kitkat1):

Can someone help me

OpenStudy (anonymous):

The answer is B 12 to the square root of 2.

OpenStudy (trojanpoem):

\[\sqrt{72} - \sqrt{8} + \sqrt{128} = \sqrt( 36 * 2) - \sqrt( 4 * 2) + \sqrt(64 * 2)\] = \[6\sqrt(2) - 2\sqrt(2) + 8 \sqrt(2) = (6 - 2 + 8) \sqrt(2) = 12\sqrt(2)\]

OpenStudy (kitkat1):

are u sure its B?

OpenStudy (trojanpoem):

See full steps above, daniel is right.

OpenStudy (kitkat1):

okay one more question Which of the following shows the correct steps to find the value of 16 to the power of 1 over 4 ? (1 point) 16 to the power of 1 over 4 equals 2 to the power of 4 to the power of 1 over 4 equals 2 to the power of 4 multiplied by 1 over 16 to the power of 1 over 4 equals 4 to the power of 4 to the power of 1 over 4 equals 4 to the power of 4 multiplied by 1 over 16 to the power of 1 over 4 equals 2 to the power of 8 to the power of 1 over 4 equals 8 to the power of 8 multiplied by 1 over 16 to the power of 1 over 4 equals 8 to the power of 2 to the power of 1 over 4 equals 2 to the power of 2 multiplied by 1 over

OpenStudy (trojanpoem):

16^0.25 = (2^4) ^0.25 = 2

OpenStudy (trojanpoem):

A

OpenStudy (kitkat1):

Simplify square root of 3 multiplied by the fifth root of 3 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 3 over 5 3 to the power of 7 over 10 Im pretty sure its D

OpenStudy (trojanpoem):

3^1/2 * 3^1/5 = 3^ (1/5 + 1/2) 1/5 + 1/2 = 2/10 + 5/10 = 7/10

OpenStudy (kitkat1):

what choice?

OpenStudy (kitkat1):

D?

OpenStudy (trojanpoem):

D

OpenStudy (trojanpoem):

I am trying to give steps, not direct answers.

OpenStudy (kitkat1):

Okay well I didnt really get them

OpenStudy (trojanpoem):

When multiplying the power is summed if the base number is the same x^ a * x^b = x^(a+b)

OpenStudy (trojanpoem):

when the question asks you to simplify an expression , with roots you find the nearest number that you can get roots for for ex sqrt(20) we know the root of 4 ? so let it be sqrt(4 * 5) = sqrt(4) * sqrt(5) ( you can split it) 2sqrt(5)

OpenStudy (trojanpoem):

When you have (x^b)^a = you can re-write it as x^ba

OpenStudy (trojanpoem):

It's just simple formulas you memorize and you got every problem.

OpenStudy (kitkat1):

So its D?

OpenStudy (trojanpoem):

Yeah dude, 1/2 + 1/5

OpenStudy (kitkat1):

okay thanks

OpenStudy (trojanpoem):

I wish you weren't just highlighting the answers.

OpenStudy (kitkat1):

I did it by myself then I just wanted to make sure